Circular No. 6975

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 52P/HARRINGTON-ABELL A. Maury, Observatoire de la Côte d’Azur, found this comet to be

undergoing what is apparently a significant outburst near m_{1} = 12.2 (m_{2} = 14.47) on CCD images on July 21.1 UT (predicted magnitude ∼ 21; cf. MPC 31833). The outburst has been confirmed by other CCD astrometrists (M. Tichý and Z. Moravec at Kleť; H. Mikuž and J. Skvarc at Crni Vrh; and R. Greimel and D. D. Balam at Victoria) and by visual observers. Available total-magnitude and coma-diameter estimates (visual unless otherwise noted): July 22.01, 10.9, 2′.8 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector; comet diffuse); 22.03, 12.3, – (Tichý and Moravec, 0.57-m reflector + CCD; m_{2} = 14.2); 22.04, 11.8, 1′.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector; low altitude); 22.07, 12.9, 1′.0 assumed (Mikuž and Skvarc, 0.36-m reflector + V filter + CCD; tail ∼ 2′ long in p.a. 258^{o}); 22.47, 11.0, – (Greimel and Balam, 1.82-m reflector + CCD + V filter).

XTE J2123–058 S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), write: “CCD photometry of the optical counterpart of this x-ray transient (IAUC 6957) has been obtained at OHP with the 1.2-m telescope during July 3–18 UT. Our V -band data yield a unique period of 5.9567 ± 0.0033 hr, identical to that reported by Tomsick et al. (IAUC 6972). In addition to this modulation, we find that the mean brightness changes from night to night. Analysis of the differences between the individual curves and the average template shows that the mean brightness is modulated by 0.3-mag (peak-to-peak) with a 7.2-day (± 0.1) period, which could be related to disk precession. The V light curve, when corrected for these baseline changes, has a full amplitude of 0.8 mag and exhibits pronounced variability near maximum light; it is remarkably similar to the published light curve of 4U 2129+47 in its active state, including a ‘shoulder’ in the descending branch at phase 0.85.”

NOVA SAGITTARII 1998 Visual magnitude estimates by K. Hornoch, Lelekovice, Czech Repub-

lic: July 2.93 UT, 11.4; 19.896, 11.5; 20.886, 11.6; 21.874, 11.5.

COMET C/1998 M6 (MONTANI) Corrigendum. This comet’s name was given incorrectly on IAUC 6974.

1998 July 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green